# Agnostic Process Tomography

### v1

@@ 1 Introduction
A fundamental problem in quantum physics and quantum learning theory is to completely characterize a quantum system by learning its evolution, a task called quantum process tomography [1, 2, 3, 4, 5, 6, 7]. Quantum process tomography has numerous applications in benchmarking [1, 3, 4, 2, 8, 9, 10, 11], error mitigation [12, 13, 14], quantum metrology [15, 16], quantum control [17], and quantum cryptography [18, 19]. Unfortunately, recovering a full description of an arbitrary quantum process requires exponential resources [6].
On the other hand, when the unknown process possesses some underlying structure, e.g., for Clifford circuits [20, 21], Clifford circuits with few non-Clifford T gates [21], shallow circuits [22, 23], junta unitaries/channels [24, 25], matrix product operators [26], circuits with a bounded number of gates [27], quantum boolean functions [28, 29], low-degree unitaries/channels [30], and Pauli channels [31, 32, 33, 34, 35, 36, 37], or when one only wishes to predict some limited properties of this process [38, 39, 8, 40, 41], efficient algorithms can be developed. However, the algorithms for the former case only apply in the realizable setting: the unknown quantum process must exactly satisfy this known structure. In practice, this assumption may not hold (e.g., due to noisy access), rendering these algorithms ineffective.
@@
Ideally, once we have such an approximation, the simple representation can be used to learn about the complex process. We outline potential applications in quantum machine learning, metrology, classical simulation, and error mitigation in Section 1.2. Notably, the analogous problem for quantum states [42, 43, 44, 45, 46] does not apply to many of these settings, motivating our generalization to quantum processes.
@@
More precisely, the problem we aim to solve is: given access to an unknown quantum process $\Phi$ and a known concept class $\mathcal{C}$ of channels, output the nearest approximation of $\Phi$ from the class $\mathcal{C}$. We refer to this task as agnostic process tomography, inspired by recent developments for agnostic state tomography [42, 43, 44, 45, 46] and classical agnostic learning [47, 48]. We are primarily interested in cases when the concept class has a simple underlying structure and consists of processes that can be efficiently implemented. One may also consider improper agnostic process tomography, in which the output of the algorithm is not in the concept class $\mathcal{C}$, but we emphasize that most applications require proper learning. We define the task formally in Section 1.1.
@@
In this work, we initiate the study of agnostic learning of quantum processes and present new efficient agnostic learning algorithms for a wide variety of concept classes, including Pauli strings, Pauli channels, quantum junta channels, low-degree channels, and a class of channels produced by $\mathsf{QAC}^{0}$ circuits. We also learn the class of bounded-gate circuits, and the associated analysis reveals that when studying agnostic process tomography, one should focus on reducing the time complexity rather than the sample complexity, analogously to the state case [42]. In addition, we prove that any agnostic state tomography algorithm can be extended to perform improper agnostic process tomography of a compatible concept class when equipped with ancilla qubits. Applying this to the agnostic state tomography algorithms of [43], we immediately obtain efficient improper agnostic learning algorithms for the classes of Clifford circuits, Clifford circuits with few T gates, and circuits consisting of a tensor product of single-qubit gates. Thus, our results answer our central question in the affirmative for these classes and also partially answer an open question of [42], regarding whether one can suitably define a notion of agnostic learning for processes and prove guarantees for the class of Clifford circuits.
@@
Notice that for the concept classes we consider in this work, learning algorithms have already been developed in the realizable setting [28, 24, 25, 30, 22, 27, 21, 20]. As a whole, our results provide new insight into when these realizable algorithms can be extended, with different technical proofs, to the agnostic setting. In particular, for simpler concept classes such as Pauli strings, Pauli channels, bounded-gate circuits and $\mathsf{QAC}^{0}$ channels, our algorithms and analysis do not differ significantly from the realizable setting [28, 27, 22]. On the other hand, for juntas and low-degree quantum channels, the realizable algorithms [24, 25, 30] fail to generalize straightforwardly to the agnostic case, requiring us to develop new, robust algorithms. Specifically, for juntas, the realizable algorithms and proofs rely crucially on a quantity (namely, the influence) that only helps characterize the unknown unitary/channel if it is indeed a junta. Similarly, for low-degree unitaries, we show that a simple agnostic extension of a realizable algorithm [30] results in a drastically increased sample/time complexity compared to both our new algorithm and the original, realizable algorithm. We elaborate more on these points in Section 1.4.
@@
We define agnostic process tomography formally in Section 1.1. Then, we describe general applications of agnostic process tomography in Section 1.2 and our results in Section 1.3. The main technical tool we use is Pauli spectrum analysis of operators and superoperators [28, 24, 22, 30, 29, 40], and we detail the ideas behind our proofs in Section 1.4.
@@ 1.1 Agnostic Process Tomography
Formally, we define agnostic process tomography as follows, where we adopt the notation of [49, 44]. We also depict the definition in Figure 1.
@@
Let $01>\epsilon,\delta>$ 0 and $\alpha\geq 1$. Let $\mathcal{C}$ be a class of completely positive trace-preserving (CPTP) maps. A learning algorithm $\mathcal{A}$ is an $(\alpha,\epsilon,\delta)$-agnostic learner with respect to the class $\mathcal{C}$ of CPTP maps if, given access to an arbitrary CPTP map $\Phi$, $\mathcal{A}$ outputs a hypothesis $\Phi^{\prime}$ such that
@@
Here, we purposefully leave some aspects of the definition flexible to capture the most general setting. In particular, we do not specify the access model for the unknown process $\Phi$ (e.g., query access, statistical query access [22, 50, 51, 52], samples from the output distribution [53, 54, 55], etc.) nor the error function $\mathsf{err}$ (e.g., diamond distance, average-case distance, Frobenius distance, etc.). Moreover, for brevity, we may refer to an $(\alpha,\epsilon,\delta)$-agnostic learner as an $\alpha$-agnostic learner and similarly refer to $\mathsf{opt}(\Phi,\mathcal{C})$ as $\mathsf{opt}$ when the other arguments are clear from context.
For our earlier motivation and the applications in Section 1.2, we focus on proper agnostic learning. We also remark that even if one succeeds at agnostic process tomography of a given concept class, the hypothesis will not be accurate if the concept class does not model the unknown channel well. Namely, despite finding a hypothesis $\Phi^{\prime}$ that is close to the optimal process $\Phi^{*}$, the optimal process itself may not approximate $\Phi$ well. One may consider, e.g., running property testing algorithms [56, 28, 57, 20, 24, 25] beforehand to check if the concept class is suitable.
Note that agnostic process tomography includes conventional process tomography [1, 2, 3, 4, 5, 6, 20, 21, 22, 23, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37] as a special case, when the unknown process is guaranteed to be in $\mathcal{C}$. However, agnostic process tomography also allows one to learn quantum processes in the presence of arbitrary noise. This greatly contrasts realizable process tomography algorithms, which are usually extremely noise-sensitive and require the unknown process to exactly satisfy some known conditions. While learning under noise is one motivation, agnostic process tomography is more broadly applicable, as discussed in the next section.
@@ 1.2 Applications
As discussed previously, selecting an appropriate concept class is crucial to applying agnostic learning algorithms to practical scenarios. By carefully choosing a concept class, we highlight a wide range of problems in near-term quantum computation that can be modelled in the framework of proper agnostic process tomography. We hope that this offers fresh perspectives and tools for further advancements in these fields. For some of these applications, our proposed algorithms are already applicable, while others require new algorithms, motivating further exploration of agnostic process tomography beyond this work. We also emphasize that these applications are unique to the agnostic process tomography framework and often cannot be addressed using agnostic state tomography [42, 43] alone.
@@
We highlight a special case of the above application for approximately classically simulating complex quantum processes. For example, given an arbitrary process, finding the closest Clifford circuit allows for efficient classical simulation [58, 59]. If the process is not too far from all Clifford circuits, which can be verified via property testing [56, 57], this will offer accurate results. This holds more generally for any classically simulable class admitting an efficient agnostic process tomography algorithm. This application is depicted in Figure 2(b).
@@
Classically, the theoretical framework of agnostic learning captures the most general settings in machine learning (ML). Namely, learning the best parameters for an ML model is equivalent to agnostic learning the concept class of all functions obtained by varying parameters of this ML model. We argue that a similar equivalence holds for quantum machine learning (QML). In QML, functions are evaluated using parameterized quantum circuits [60], or more generally, quantum channels (such as a Linear Combinations of Unitaries (LCU) [61, 62, 63]). By choosing an appropriate access model and error, agnostic process tomography of the class of channels expressible by an ansatz allows one to find the best parameters for the QML model. In this sense, efficient agnostic process tomography of a model class implies efficient trainability. Thus, studying agnostic process tomography under various access models may lead to new insights into QML for both quantum and classical problems. For example, many QML algorithms learn from expectation values of observables on output states of parameterized quantum circuits, which can be modelled as quantum statistical queries[64, 65, 50, 51, 22, 55]. Thus, agnostic process tomography from quantum statistical queries could be a promising direction for new insights in QML. This application is depicted in Figure 2(c).
@@
Quantum metrology is considered one of the most significant real-world applications of quantum technology [16]. Here, one aims to estimate an unknown parameter $\theta$ with precision limited by the fundamental bounds of measurement [15], given access to a known class of parameterized quantum channels $\{\mathcal{E}_{\theta}\}_{\theta}$ called the encoding. By probing the device modelled by the channel and applying general quantum measurements to the outputs, one can construct an estimator $\hat{\theta}$ of the unknown parameter with desirable statistical properties. However, in practice, the encoding may be corrupted by unknown environmental noise. Due to the challenges of characterizing noise and its effect on parameter estimation, noisy quantum metrology is less well-studied, and results are restricted to specific classes of noise [66, 67, 68, 69, 70, 71, 72]. Many standard protocols are also highly noise-sensitive [16, 73, 74, 75]. Considering the known class $\{\mathcal{E}_{\theta}\}_{\theta}$ as our concept class, one can model the problem of extracting a desirable estimator as an agnostic process tomography problem, which may provide us with new insights into metrology in noisy settings. While noiseless metrology is close to conventional tomography, the parameter-encoding channel becomes unknown when transformed by realistic noise, which is more suitable for agnostic learning. This perspective is also applicable in adversarial scenarios, where an unknown attack channel affects the encoding [76, 77]. We are unsure how this approach would compare to the standard quantum limit (the Heisenberg limit is unattainable for general noise without quantum error correction [78]) but nevertheless believe this is interesting to explore. This application is depicted in Figure 2(c).
@@
Error mitigation protocols are often based on machine learning [79, 80] or require a strong characterization of the noise channel beforehand [12, 81]. We argue that agnostic process tomography can be used to generalize a machine-learning-based error mitigation protocol called Clifford Data Regression [79]. In this setting, we wish to mitigate a fixed, arbitrary noise channel $\mathcal{E}$ on a quantum device. We assume that $\mathcal{E}$ is well-approximated by a parameterized class of noise channels $\{\mathcal{E}_{\theta}\}_{\theta}$ such that, when $\theta$ is known, the errors can be mitigated via classical post-processing. This is relevant when methods that map arbitrary noise to a known class, e.g., Pauli twirling [81, 82, 83, 84], are unavailable. We propose the following protocol: 1) Execute various classically-simulable circuits on the quantum device to obtain noisy data. 2) Simulate the same circuits classically to obtain ideal, noiseless data. 3) Using the noisy and ideal data, perform agnostic process tomography with respect to $\{\mathcal{E}_{\theta}\}_{\theta}$. This procedure outputs an approximation $\mathcal{E}_{\theta^{\prime}}$ of the best representative of the arbitrary noise channel $\mathcal{E}$ from $\{\mathcal{E}_{\theta}\}_{\theta}$. Then, when implementing new circuits on this device, we can use $\mathcal{E}_{\theta^{\prime}}$ to mitigate noise via classical post-processing. Our new protocol can be seen as providing the pre-characterization of error often needed for error mitigation protocols such as probabilistic error cancellation [81]. Note that in this setting, our access to the unknown process is through the pairs of ideal and noisy data rather than query access.
We consider the class of Pauli channels $\sum_{x}p_{x}\sigma_{x}(\cdot)\sigma_{x}$, parameterized by an underlying probability distribution $\{p_{x}\}_{x}$, as an example of our protocol for error mitigation. Our agnostic tomography algorithm for Pauli channels in Section 5.3 only requires data obtained by performing measurements in the basis of Choi states of the Pauli operators. The approximated noise can then be used to mitigate errors in new data using standard techniques for Pauli channel noise [81, 85]. Note a limitation is that our agnostic learning algorithm for Pauli channels requires the use of ancilla qubits. This example also indicates that for this protocol to work, agnostic process tomography cannot be used in a black box manner. Instead, the agnostic algorithm should dictate what data to gather from the noisy quantum device. While we do not show an immediate application, we believe that this approach can result in new protocols for error mitigation for specific classes of noise. This application is depicted in Figure 2(d).

### v2

@@ 1 Introduction
A fundamental problem in quantum physics and quantum learning theory is to completely characterize a quantum system by learning its evolution, a task called quantum process tomography [mohseni2008quantum, o2004quantum, scott2008optimizing, chuang1997prescription, d2001quantum, haah2023query, poyatos1997complete]. Quantum process tomography has numerous applications in benchmarking [mohseni2008quantum, scott2008optimizing, chuang1997prescription, o2004quantum, levy2024classical, huang2022foundations, merkel2013self, blume2017demonstration], error mitigation [endo2021hybrid, seif2023shadow, cai2023quantum], quantum metrology [giovannetti2006quantum, giovannetti2011advances], quantum control [kim2014quantum], and quantum cryptography [arapinis2021quantum, schuster2024random]. Unfortunately, recovering a full description of an arbitrary quantum process requires exponential resources [haah2023query].
On the other hand, when the unknown process possesses some underlying structure, e.g., for Clifford circuits [low2009learning, lai2022learning], Clifford circuits with few non-Clifford T gates [lai2022learning], shallow circuits [nadimpalli2024pauli, huang2024learning], junta unitaries/channels [chen2023testing, bao2023testing], matrix product operators [torlai2023quantum], circuits with a bounded number of gates [zhao2023learning], quantum boolean functions [montanaro2008quantum, rouze2024quantum], low-degree unitaries/channels [arunachalam2024learning], and Pauli channels [harper2020efficient, harper2021fast, flammia2020efficient, chen2022quantum, flammia2021pauli, fawzi2023lower, chen2024tight], or when one only wishes to predict some limited properties of this process [kunjummen2023shadow, huang2023learning, levy2024classical, chen2024predicting, chung2018sample], efficient algorithms can be developed. However, the algorithms for the former case only apply in the realizable setting: the unknown quantum process must exactly satisfy this known structure. In practice, this assumption may not hold (e.g., due to noisy access), rendering these algorithms ineffective.
@@
Ideally, once we have such an approximation, the simple representation can be used to learn about the complex process. We outline potential applications in quantum machine learning, metrology, classical simulation, and error mitigation in Section 1.2. Notably, the analogous problem for quantum states [grewal2024agnostic, chen2024stabilizer, anshu2024survey, badescu2021improved, grewal2024improved] does not apply to many of these settings, motivating our generalization to quantum processes.
@@
More precisely, the problem we aim to solve is: given access to an unknown quantum process $\Phi$ and a known concept class $\mathcal{C}$ of channels, output the nearest approximation of $\Phi$ from the class $\mathcal{C}$. We refer to this task as agnostic process tomography, inspired by recent developments for agnostic state tomography [grewal2024agnostic, chen2024stabilizer, anshu2024survey, badescu2021improved, grewal2024improved] and classical agnostic learning [kearns1992toward, haussler1992decision]. We are primarily interested in cases when the concept class has a simple underlying structure and consists of processes that can be efficiently implemented. One may also consider improper agnostic process tomography, in which the output of the algorithm is not in the concept class $\mathcal{C}$, but we emphasize that most applications require proper learning. We define the task formally in Section 1.1.
@@
In this work, we initiate the study of agnostic learning of quantum processes and present new efficient agnostic learning algorithms for a wide variety of concept classes, including Pauli strings, Pauli channels, quantum junta channels, low-degree channels, and a class of channels produced by $\mathsf{QAC}^{0}$ circuits. We also learn the class of bounded-gate circuits, and the associated analysis reveals that when studying agnostic process tomography, one should focus on reducing the time complexity rather than the sample complexity, analogously to the state case [grewal2024agnostic]. In addition, we prove that any agnostic state tomography algorithm can be extended to perform improper agnostic process tomography of a compatible concept class when equipped with ancilla qubits. Applying this to the agnostic state tomography algorithms of [chen2024stabilizer], we immediately obtain efficient improper agnostic learning algorithms for the classes of Clifford circuits, Clifford circuits with few T gates, and circuits consisting of a tensor product of single-qubit gates. Thus, our results answer our central question in the affirmative for these classes and also partially answer an open question of [grewal2024agnostic], regarding whether one can suitably define a notion of agnostic learning for processes and prove guarantees for the class of Clifford circuits.
@@
Notice that for the concept classes we consider in this work, learning algorithms have already been developed in the realizable setting [montanaro2008quantum, chen2023testing, bao2023testing, arunachalam2024learning, nadimpalli2024pauli, zhao2023learning, lai2022learning, low2009learning]. As a whole, our results provide new insight into when these realizable algorithms can be extended, with different technical proofs, to the agnostic setting. In particular, for simpler concept classes such as Pauli strings, Pauli channels, bounded-gate circuits and $\mathsf{QAC}^{0}$ channels, our algorithms and analysis do not differ significantly from the realizable setting [montanaro2008quantum, zhao2023learning, nadimpalli2024pauli]. On the other hand, for juntas and low-degree quantum channels, the realizable algorithms [chen2023testing, bao2023testing, arunachalam2024learning] fail to generalize straightforwardly to the agnostic case, requiring us to develop new, robust algorithms. Specifically, for juntas, the realizable algorithms and proofs rely crucially on a quantity (namely, the influence) that only helps characterize the unknown unitary/channel if it is indeed a junta. Similarly, for low-degree unitaries, we show that a simple agnostic extension of a realizable algorithm [arunachalam2024learning] results in a drastically increased sample/time complexity compared to both our new algorithm and the original, realizable algorithm. We elaborate more on these points in Section 1.4.
@@
We define agnostic process tomography formally in Section 1.1. Then, we describe general applications of agnostic process tomography in Section 1.2 and our results in Section 1.3. The main technical tool we use is Pauli spectrum analysis of operators and superoperators [montanaro2008quantum, chen2023testing, nadimpalli2024pauli, arunachalam2024learning, rouze2024quantum, chen2024predicting], and we detail the ideas behind our proofs in Section 1.4.
@@ 1.1 Agnostic Process Tomography
Formally, we define agnostic process tomography as follows, where we adopt the notation of [caro2023classical, anshu2024survey]. We also depict the definition in Figure 1.
@@
Let $1>\epsilon,\delta>0$ and $\alpha\geq 1$. Let $\mathcal{C}$ be a class of completely positive trace-preserving (CPTP) maps. A learning algorithm $\mathcal{A}$ is an $(\alpha,\epsilon,\delta)$-agnostic learner with respect to the class $\mathcal{C}$ of CPTP maps if, given access to an arbitrary CPTP map $\Phi$, $\mathcal{A}$ outputs a hypothesis $\Phi^{\prime}$ such that
@@
Here, we purposefully leave some aspects of the definition flexible to capture the most general setting. In particular, we do not specify the access model for the unknown process $\Phi$ (e.g., query access, statistical query access [nadimpalli2024pauli, angrisani2023learning, wadhwa2023learning, wadhwa2024noise], samples from the output distribution [hinsche2023one, nietner2023average, nietner2023unifying], etc.) nor the error function $\mathsf{err}$ (e.g., diamond distance, average-case distance, Frobenius distance, etc.). Moreover, for brevity, we may refer to an $(\alpha,\epsilon,\delta)$-agnostic learner as an $\alpha$-agnostic learner and similarly refer to $\mathsf{opt}(\Phi,\mathcal{C})$ as $\mathsf{opt}$ when the other arguments are clear from context.
For our earlier motivation and the applications in Section 1.2, we focus on proper agnostic learning. We also remark that even if one succeeds at agnostic process tomography of a given concept class, the hypothesis will not be accurate if the concept class does not model the unknown channel well. Namely, despite finding a hypothesis $\Phi^{\prime}$ that is close to the optimal process $\Phi^{*}$, the optimal process itself may not approximate $\Phi$ well. One may consider, e.g., running property testing algorithms [montanaro2013survey, montanaro2008quantum, wang2011property, low2009learning, chen2023testing, bao2023testing] beforehand to check if the concept class is suitable.
Note that agnostic process tomography includes conventional process tomography [mohseni2008quantum, o2004quantum, scott2008optimizing, chuang1997prescription, d2001quantum, haah2023query, low2009learning, lai2022learning, nadimpalli2024pauli, huang2024learning, torlai2023quantum, zhao2023learning, montanaro2008quantum, rouze2024quantum, arunachalam2024learning, harper2020efficient, harper2021fast, flammia2020efficient, chen2022quantum, flammia2021pauli, fawzi2023lower, chen2024tight] as a special case, when the unknown process is guaranteed to be in $\mathcal{C}$. However, agnostic process tomography also allows one to learn quantum processes in the presence of arbitrary noise. This greatly contrasts realizable process tomography algorithms, which are usually extremely noise-sensitive and require the unknown process to exactly satisfy some known conditions. While learning under noise is one motivation, agnostic process tomography is more broadly applicable, as discussed in the next section.
@@ 1.2 Applications
As discussed previously, selecting an appropriate concept class is crucial to applying agnostic learning algorithms to practical scenarios. By carefully choosing a concept class, we highlight a wide range of problems in near-term quantum computation that can be modelled in the framework of proper agnostic process tomography. We hope that this offers fresh perspectives and tools for further advancements in these fields. For some of these applications, our proposed algorithms are already applicable, while others require new algorithms, motivating further exploration of agnostic process tomography beyond this work. We also emphasize that these applications are unique to the agnostic process tomography framework and often cannot be addressed using agnostic state tomography [grewal2024agnostic, chen2024stabilizer] alone.
@@
We highlight a special case of the above application for approximately classically simulating complex quantum processes. For example, given an arbitrary process, finding the closest Clifford circuit allows for efficient classical simulation [aaronson2004improved, gottesman1998theory]. If the process is not too far from all Clifford circuits, which can be verified via property testing [montanaro2013survey, wang2011property], this will offer accurate results. This holds more generally for any classically simulable class admitting an efficient agnostic process tomography algorithm. This application is depicted in Figure 2(b).
@@
Classically, the theoretical framework of agnostic learning captures the most general settings in machine learning (ML). Namely, learning the best parameters for an ML model is equivalent to agnostic learning the concept class of all functions obtained by varying parameters of this ML model. We argue that a similar equivalence holds for quantum machine learning (QML). In QML, functions are evaluated using parameterized quantum circuits [cerezo2021variational], or more generally, quantum channels (such as a Linear Combinations of Unitaries (LCU) [childs2012hamiltonian, heredge2024non, coyle2024training]). By choosing an appropriate access model and error, agnostic process tomography of the class of channels expressible by an ansatz allows one to find the best parameters for the QML model. In this sense, efficient agnostic process tomography of a model class implies efficient trainability. Thus, studying agnostic process tomography under various access models may lead to new insights into QML for both quantum and classical problems. For example, many QML algorithms learn from expectation values of observables on output states of parameterized quantum circuits, which can be modelled as quantum statistical queries[arunachalam2020quantum, arunachalam2024role, angrisani2023learning, wadhwa2023learning, nadimpalli2024pauli, nietner2023unifying]. Thus, agnostic process tomography from quantum statistical queries could be a promising direction for new insights in QML. This application is depicted in Figure 2(c).
@@
Quantum metrology is considered one of the most significant real-world applications of quantum technology [giovannetti2011advances]. Here, one aims to estimate an unknown parameter $\theta$ with precision limited by the fundamental bounds of measurement [giovannetti2006quantum], given access to a known class of parameterized quantum channels $\{\mathcal{E}_{\theta}\}_{\theta}$ called the encoding. By probing the device modelled by the channel and applying general quantum measurements to the outputs, one can construct an estimator $\hat{\theta}$ of the unknown parameter with desirable statistical properties. However, in practice, the encoding may be corrupted by unknown environmental noise. Due to the challenges of characterizing noise and its effect on parameter estimation, noisy quantum metrology is less well-studied, and results are restricted to specific classes of noise [gorecki2022quantum, demkowicz2012elusive, escher2011general, ji2008parameter, zhou2018achieving, demkowicz2017adaptive, zhou2024achieving]. Many standard protocols are also highly noise-sensitive [giovannetti2011advances, ono2010effects, gilbert2008use, rubin2007loss]. Considering the known class $\{\mathcal{E}_{\theta}\}_{\theta}$ as our concept class, one can model the problem of extracting a desirable estimator as an agnostic process tomography problem, which may provide us with new insights into metrology in noisy settings. While noiseless metrology is close to conventional tomography, the parameter-encoding channel becomes unknown when transformed by realistic noise, which is more suitable for agnostic learning. This perspective is also applicable in adversarial scenarios, where an unknown attack channel affects the encoding [shettell2022cryptographic, huang2019cryptographic]. We are unsure how this approach would compare to the standard quantum limit (the Heisenberg limit is unattainable for general noise without quantum error correction [zhou2024limits]) but nevertheless believe this is interesting to explore. This application is depicted in Figure 2(c).
@@
Error mitigation protocols are often based on machine learning [czarnik2021error, strikis2021learning] or require a strong characterization of the noise channel beforehand [endo2021hybrid, van2023probabilistic]. We argue that agnostic process tomography can be used to generalize a machine-learning-based error mitigation protocol called Clifford Data Regression [czarnik2021error]. In this setting, we wish to mitigate a fixed, arbitrary noise channel $\mathcal{E}$ on a quantum device. We assume that $\mathcal{E}$ is well-approximated by a parameterized class of noise channels $\{\mathcal{E}_{\theta}\}_{\theta}$ such that, when $\theta$ is known, the errors can be mitigated via classical post-processing. This is relevant when methods that map arbitrary noise to a known class, e.g., Pauli twirling [van2023probabilistic, gupta2024probabilistic, geller2013efficient, wallman2016noise], are unavailable. We propose the following protocol: 1) Execute various classically-simulable circuits on the quantum device to obtain noisy data. 2) Simulate the same circuits classically to obtain ideal, noiseless data. 3) Using the noisy and ideal data, perform agnostic process tomography with respect to $\{\mathcal{E}_{\theta}\}_{\theta}$. This procedure outputs an approximation $\mathcal{E}_{\theta^{\prime}}$ of the best representative of the arbitrary noise channel $\mathcal{E}$ from $\{\mathcal{E}_{\theta}\}_{\theta}$. Then, when implementing new circuits on this device, we can use $\mathcal{E}_{\theta^{\prime}}$ to mitigate noise via classical post-processing. Our new protocol can be seen as providing the pre-characterization of error often needed for error mitigation protocols such as probabilistic error cancellation [van2023probabilistic]. Note that in this setting, our access to the unknown process is through the pairs of ideal and noisy data rather than query access.
We consider the class of Pauli channels $\sum_{x}p_{x}\sigma_{x}(\cdot)\sigma_{x}$, parameterized by an underlying probability distribution $\{p_{x}\}_{x}$, as an example of our protocol for error mitigation. Our agnostic tomography algorithm for Pauli channels in Section 5.3 only requires data obtained by performing measurements in the basis of Choi states of the Pauli operators. The approximated noise can then be used to mitigate errors in new data using standard techniques for Pauli channel noise [van2023probabilistic, ferracin2024efficiently]. Note a limitation is that our agnostic learning algorithm for Pauli channels requires the use of ancilla qubits. This example also indicates that for this protocol to work, agnostic process tomography cannot be used in a black box manner. Instead, the agnostic algorithm should dictate what data to gather from the noisy quantum device. While we do not show an immediate application, we believe that this approach can result in new protocols for error mitigation for specific classes of noise. This application is depicted in Figure 2(d).